### v2

Abstract
Free-text responses are commonly collected in psychological studies, providing rich qualitative insights that quantitative measures may not capture. Labeling curated topics of research interest in free-text data by multiple trained human coders is typically labor-intensive and time-consuming. Though large language models (LLMs) excel in language processing, LLM-assisted labeling techniques relying on closed-source LLMs cannot be directly applied to free-text data, without explicit consent for external use.
In this study, we propose a framework of assembling locally-deployable LLMs to enhance the labeling of predetermined topics in free-text data under privacy constraints. Analogous to annotation by multiple human raters, this framework leverages the heterogeneity of diverse open-source LLMs. The ensemble approach seeks a balance between the agreement and disagreement across LLMs, guided by a relevancy scoring methodology that utilizes embedding distances between topic descriptions and LLMs’ reasoning. We evaluated the ensemble approach using both publicly accessible Reddit data from eating disorder related forums, and free-text responses from eating disorder patients, both complemented by human annotations.
We found that: (1) there is heterogeneity in the performance of labeling among same-sized LLMs, with some showing low sensitivity but high precision, while others exhibit high sensitivity but low precision. (2) Compared to individual LLMs, the ensemble of LLMs achieved the highest accuracy and optimal precision-sensitivity trade-off in predicting human annotations. (3) The relevancy scores across LLMs showed greater agreement than dichotomous labels, indicating that the relevancy scoring method effectively mitigates the heterogeneity in LLMs’ labeling.
keywords:
Free-text Data, Ensemble of LLMs, Open-Source LLMs, Topic Annotation, Relevancy Scoring
1 Introduction
In psychology, free-form text responses and qualitative data have been effectively collected since the beginning of scientific inquiry [1, 2]. To perform quantitative analyses on psychological textual data, curated topics such as symptom manifestations [3], psychosis risk factors [4], social risk factors [5], and pathological behaviors [6] must first be labeled (coded) and tailored to specific research interests. Conventionally, this process involves multiple human coders independently labeling and rating topics of interest, with reliability evaluated through the level of agreement [7]. However, this process is typically time-consuming and labor-intensive, involving both the careful training of human coders as well as their efforts in reading and labeling large volumes of text.
Large language models (LLMs) are pre-trained statistical language models renowned for their ability to process and generate human-like language [8]. Research have shown that larger-sized LLMs, such as the GPTs [9, 10, 11], offer significant opportunities for facilitating data annotation [12, 13, 14, 15, 16]. However, ethical concerns regarding privacy and safety of free-text data should be central to the development of LLM-assisted methodologies [2, 17, 18]. Without explicit consent for external use, closed-source LLMs, or large-sized LLMs that cannot be deployed locally, are often unsuitable for direct application to free-text data collected in clinical or research settings [17, 18, 19].
In this study, we focused on enhancing the accuracy and reliability of labeling curated topics on free-text data, using ensembles of small-sized (7-8 billion parameters), and locally-deployable LLMs, considering both security and computational constraints. Instead of inferring mental health conditions or psychological constructs, our work focused on labeling specific topics that require detailed descriptions tailored to particular research interests. The proposed ensembling framework was not designed to replace human annotation but rather to alleviate the labor burden by providing human raters with a preliminary label and relevancy score for each text.
The rest of the article is organized as follows: We dedicate the remainder of the introduction (Section 1) to detailing the background of LLMs and their applications in data annotation generally, and specifically in psychology (Section 1.1). Then we discuss the rationale behind using ensembles of locally-deployable LLMs for topic labeling on free-text data (Section 1.2), followed by a related work that utilized LLM ensembles for data annotation (Section 1.3). In the methods section (Section 2), we introduce our ensemble framework (Section 2.1), and detail the evaluations using both a public and private psychological dataset (Section 2.2). Section 3 compares the ensemble with individual LLMs in terms of fidelity to human labeling. Section 4 contains a discussion of our new methodology.
1.1 LLMs for data annotation
Large language models are large-scale, pre-trained statistical language models known for their capacity and versatility in processing and generating human-like language [8]. LLMs can be categorized as open-source or closed-source based on developers’ policies [2]. Open-source models like LLaMA [20] support local deployment, enabling dataset customization and ensuring data privacy [8]. In contrast, closed-source models, such as commercial GPTs [9, 10, 11], are controlled by third parties and accessible via application programming interfaces (APIs). To adapt LLMs to specific tasks in specialized domains, two methodologies are commonly used. Fine-tuning involves further training a pre-trained model on a task-specific dataset to enhance its capabilities [21]. Prompting techniques, in contrast, adapt LLMs without additional training. The main techniques are zero-shot learning, which uses crafted prompts to direct the model’s output, and few-shot learning, which extends this by providing a few examples within the prompt, helping the model to infer the task [9].
Research has shown that advanced LLMs, exemplified by the GPTs, present promising opportunities for facilitating data annotation [12, 13, 14, 15, 16]. The use of larger-sized and closed-source LLMs typically outperform smaller-sized and open-source LLMs. Gilardi et al. [14] demonstrated that ChatGPT outperforms open-source LLMs in generic annotation tasks, including relevance, stance, topic, and frame detection. Alizadeh et al. [15] demonstrated that fine-tuning can enhance the performance of open-source LLMs in text classification tasks within political science, getting closer to the performance of zero-shot prompting with GPT-3.5 and GPT-4. Chew et al. [16] proposed an LLM-assisted content analysis approach using GPT-3.5, which involves iterative steps by both humans and the LLM: codebook development, labeling, and agreement evaluation, repeated until a satisfactory level of human-LLM agreement is achieved. They found that GPT-3.5 can often perform labeling at agreement levels comparable to those of human coders.
In psychology, besides data augmentation [22, 23], resource enrichment [24], and developing new LLM agents for psychological applications through fine-tuning or prompting [25, 26], LLMs have been employed as classification models to label the presence or absence of mental health conditions, such as suicide and depression [27, 28], or identifying more detailed information about a mental health condition, such as severity levels and subtypes [29, 30, 2]. The usage of closed-source LLMs such as GPTs are at the forefront of these applications [2, 31]. Additionally, LLMs have been used to annotate psychological constructs. Rathje et al. [32] demonstrated that GPT-3.5-Turbo, GPT-4, and GPT-4-Turbo can accurately detect various psychological constructs (sentiment, discrete emotions, offensiveness, and moral foundations) across multiple languages using simple prompts and no additional training data, as validated by manual annotators. Peters and Matz [33] indicated that GPT-3.5 and GPT-4 can infer personality traits and psychological dispositions from social media users. Kjell et al. [19] provided evidence that, with careful validation of targeted deployment scenarios, the latest LLM technologies can transition psychological assessment from traditional rating scales to natural language communication.
However, several considerations must be addressed before using LLMs to label curated psychological topics in free-text data collected from clinical or research settings. First, as a comprehensive review by Hua et al. [2] pointed out, labels of mental health conditions in existing datasets are typically not curated, with substantial variability in their definitions across studies. Curated psychological topics tailored to individual research needs—such as pathological behaviors, social or cultural factors related to a mental condition, and potential risk factors hypothesized by researchers—should be differentiated from clinical diagnoses of mental health conditions (e.g., depression or suicide) and the assessment of abstract constructs (e.g., sentiment or personality traits). Given the diversity of such curated topics and their study-specific definitions, evaluating LLMs’ performance demands a higher level of precision, complemented with human annotations tailored to each topic’s definition. In this study, we focused on curated topics with detailed definitions, ensuring that human annotations in our empirical evaluations closely aligned with the LLMs’ labeling process.
Second, ethical concerns regarding the privacy and safety of free-text data are pivotal in developing LLM-assisted methodologies [2, 17, 18]. Current applications frequently utilize closed-source, large-scale LLMs, such as GPTs. While these models provide strong language processing capabilities, they require data to be sent beyond the researchers’ control. Sharing protected health-related data (such as clinician notes) via APIs and storing it in non-private environments, whether temporarily or long-term, poses risks to data privacy and safety. Moreover, without explicit consent for external use—which is rarely obtained if at all—closed-source LLMs are unsuitable for direct application to free-text data collected in clinical or research settings [17, 18, 19]. Larger-sized LLMs that can hardly be deployed locally in a secure computing environment, such as the open-source Llama-3.1-405B [34], may also be unsuitable for handling protected free-text data. Thus, our work focuses on improving the accuracy and reliability of topic labeling by assembling diverse, open-source, small-sized (7-8 billion parameters), and locally-deployable LLMs, under security and computational constraints.
Lastly, the use of a single LLM in data annotation tasks is commonplace in current applications. However, studies have shown that diverse LLMs exhibit varying capacities across tasks [35, 36], and it is reasonable to expect heterogeneity in their performance when annotating diverse topics with granular definitions. As highlighted in a survey on LLM-assisted data annotation [12], employing the Mixture of Experts concept [37] and ensembling multiple LLMs can leverage this heterogeneity to enhance overall performance and computational efficiency [38], which we elaborate on in the next section.
1.2 Ensemble of LLMs for curated topic labeling on free-text data
The usage of multiple LLMs is analogous to multiple human raters who are limited to their own knowledge and biases. Studies have shown that diverse LLMs exhibit varying capacities in performing inference tasks such as coding, taxonomic labeling, and math [35, 36, satpute2024can]. We expect to observe heterogeneity in the performances of diverse open-source, same-sized LLMs when labeling various curated psychological topics. Besides differences in parameter sizes and model architectures [39], the heterogeneity in performance among LLMs also stems from the varied training datasets and preprocessing steps each LLM was exposed to, or inherited if the LLM is a variant of another [40].
The main objective of ensembling diverse LLMs is to leverage the heterogeneity in LLMs and achieve reliable performance in topic labeling tasks. Ensemble learning of LLMs is typically performed either by ensembling model weights or by combining diverse outputs [41, 42]. A popular technique–LLM routing–involves training a router model using various benchmark datasets to assign LLMs to their specialized tasks and applications, aiming to balance performance and cost [43, 42, 44, 45, 46, 47]. However, this technique is not directly applicable to our task of curated topic labeling for free-text data. (1) Current router models require the inclusion of large-sized, closed-source LLMs such as GPT at the high-cost, high-performance end of the performance-cost trade-off [42, 44, 46], which is not directly applicable to free-text data. (2) The development of router models has primarily focused on assigning LLMs to their specialties in generic tasks, such as commonsense reasoning, conversation, mathematics, and code generation. Routing LLMs for the specific task of topic labeling has not been investigated, except for one study that will be discussed in detail in the related work section. The effectiveness of using this generic technique to label curated psychological topics remains unestablished. (3) Training a router model requires a pre-specified collection of LLMs and extensive training datasets, making it challenging to scale to new LLM selections, as well as diverse datasets and research needs in psychology studies.
Therefore, in this study, we propose an ensembling framework that is tailored to the task of curated topic labeling for free-text data. This framework is scalable in LLM selections and can be readily adapted to individual research needs without requiring the training of a new model. The framework leverages the heterogeneity of diverse, open-source, and locally-deployable LLMs to enhance the accuracy and reliability of determining the presence or absence of a topic in a given text. The ensembling seeks a balance between the agreement and disagreement across multiple LLMs, guided by a novel relevancy scoring methodology. We evaluated the ensembling framework using both a public dataset related to a mental health condition and a protected clinical dataset, both complemented by human annotations. For the public dataset, we also employed the large-sized GPT-4o as a judge [48]. This framework was not designed to replace human evaluation but rather to alleviate the labor burden by providing human raters with a preliminary label and relevancy score for each text.
1.3 Related work
A very recent work utilized the ensemble of LLMs for data annotation tasks [49]. Farr et al. [49] proposed an ensembling methodology that aligns multiple LLMs in a chain, routing subsets of data to subsequent models based on the classification uncertainty measured for the current model. This ensemble is an extension of the LLM routing technique but does not require to train a router model. It was shown to outperform individual models in the chain and reduce the use of high-cost LLMs. However, our ensembling framework differs significantly from theirs. First, as an extension of the LLM routing technique, their designs incorporated both open-source smaller-sized LLMs and a large-sized, closed-source GPT-4o. The extent to which performance improvements attributed to the use of GPT-4o remains unclear. Given the constraints of free-text data, our method mainly focuses on enhancing the performance of open-source deployable LLMs through ensembling. Second, their method suggests that more robust models be placed later in the chain, with the robustness indicated by a progression from smaller to larger model sizes. In our study, we aim to evaluate the advancements of ensembling compared to individual LLMs, controlling for the same model sizes. In addition to the F1-score assessed in their experiment, we evaluated the performance through multiple lenses, including precision, sensitivity and inter-rater reliability. Lastly, the labeling tasks evaluated in their empirical experiments—stance detection, misinformation detection, and ideology detection—are abstract components of text, that do not directly align with the needs for precise labeling of curated topics in psychological studies.
2 Methods
2.1 Ensemble LLMs for Topic Labeling
In this section, we present the detailed methodologies for the four stages of the ensemble framework illustrated in Figure 1: (1) topic labeling by each selected open-source LLM; (2) relevancy scoring based on the embedding distance between each LLM’s reasoning and the topic description; (3) agreement evaluation among LLMs and the removal of potential outliers; and (4) the ensemble of scores and labels from the final set of LLMs.
Figure 1: The ensemble framework is illustrated in two parts. The upper panel depicts the pipeline for topic labeling and relevancy scoring. Each LLM is prompted with text and a detailed topic description, and returns both a dichotomous label and related phrases identified from the text. Relevancy scores are calculated using the cosine similarity between the embeddings of the topic description and the returned phrases. The lower panel represents the ensemble methodology, which leverages the heterogeneous performances of diverse open-source, deployable-sized language models. The final ensemble of labels is determined by removing false positives from the union of positive cases, using the optimal threshold on the ensemble of relevancy scores that predicts the intersection of positive cases with the highest F1-score.
2.1.1 Topic Labeling
A collection of open-source language models should be used to conduct the same labeling and scoring task for a topic of interest. In this experimentation, we selected four small-sized (7-8 billion parameters) language models that are open-source on huggingface: Llama-3.1-8B-Instruct [34], Qwen2.5-7B-Instruct [50], Mistral-7B-Instruct-v0.3 [51], and Vicuna-7b-v1.5 [48, 52]. Each model can be deployed in a secure computing environment and is recognized for its capacity in text generation inference [53]. To determine the presence of a topic in a given textual data, we prompted each LLM with the text and a detailed description of the topic. The prompt design is shown in the upper panel of Figure 1. Although techniques such as few-shot prompting [54], chain-of-thought [55], and other prompt engineering techniques [56, 57] can be employed, this study focuses on comparing the ensemble performance with individual performances of LLMs in topic labeling tasks. We used the same prompt scheme across all LLMs, allowing diverse models to reveal their heterogeneous capacities in performing the same task.
2.1.2 Relevancy Scoring
Recent transformer-based embeddings [58], exemplified by BERT (Bidirectional Encoder Representations from Transformers) models [59], mark a significant advancement in capturing contextualized word representations by simultaneously enlabeling words and their relationships within a sentence as numerical embedding vectors [60]. The distance between two sentences can be quantified using the cosine similarity between their vectors in the embedding space [61].
In our prompt design, each LLM is asked to return both a dichotomous label indicating the presence of a given topic in the text, as well as the topic-related phrases identified within the text, which reflect the LLM’s reasoning. Subsequently, a relevancy score is calculated using the cosine similarity between the embeddings of the topic description and the related phrases returned by an LLM. This score measures the extent to which the evidence provided by an LLM is relevant to the queried topic and reflects how explicitly the text discusses the topic.
In our experiments, embeddings are generated using the pre-trained Sentence-BERT model all-mpnet-base-v2 [60], which maps sentences (truncated to a maximum of 384 words) into a 768-dimensional vector space. This model, trained on over 1 billion sentences, has demonstrated optimal performance in semantic search according to the massive text embedding benchmark leaderboard [62].
Lastly, for a given topic description, we calculated a baseline cosine similarity score between the description and an empty string. We set the observed scores under the baseline to zero, and subtracted the baseline score from the observed score, yielding a final relevancy score ranging from 0 to 1.
2.1.3 Agreement Evaluation
Before ensembling, we conducted an agreement evaluation on the labels and scores returned by the LLMs. Gwet’s AC1 [63] and Fleiss’s Kappa [64] are the preferred inter-rater reliability metrics in scenarios where the topic of interest is rarely prevalent [16, 64]. Other metrics, such as Cohen’s Kappa [65], may result in the “high agreement, low reliability” paradox in rare event labeling [66, 64, 67]. To evaluate the agreement among the continuous relevancy scores across LLMs, we categorized the scores into ten ordinal levels, each spanning 0.1, before calculating unweighted Gwet’s AC1 and Fleiss’s Kappa.
Although the ensemble of LLMs should benefit from the heterogeneity of diverse LLMs, those whose labels and scores significantly deviate from the rest may detrimentally affect the ensembling process. Thus, to remove potential outlier LLMs, we calculated the increase in Gwet’s AC1 when each LLM was excluded from the full collection. If this increase surpasses a predetermined threshold, the LLM should be excluded from the ensembling candidates. The choice of threshold depends on the needs of specific applications, and in our empirical evaluation, we required the increase to surpass 10% of the overall AC1.
2.1.4 Ensembling
From each LLM, we retrieve a binary label and a relevancy score ranging from 0 to 1 if the label is positive. An ensemble of relevancy scores can be generated using the principal component that explains the most variability in the data, through principal component analysis (PCA), provided that the sample size of the text data is sufficient. Then, the PCA ensemble of scores is used to inform the ensembling of labels.
The primary objective in ensembling the labels is to eliminate potential false positives from the union of positive labels returned by the LLMs. The union of positive labels from multiple LLMs inherits all the false positives from each LLM. Conversely, the intersection of positive labels, defined as those agreed upon by more than half of the LLMs, could be too conservative, potentially failing to include all true positives. Thus, the ensembling of labels involves finding a balance between the union and the intersection, guided by the PCA ensemble of relevancy scores. Specifically, we examined each threshold on the scale of the PCA ensemble of scores and identified the optimal threshold that predicts the intersection of positive cases with the highest F1-score. This threshold represents the optimized trade-off between sensitivity and precision, particularly for topics with rare prevalence [68]. Subsequently, we excluded those cases in the union of positive cases whose scores fell below this threshold as false positives.
2.2 Empirical evaluation
Eating disorders (ED) are serious psychiatric illnesses that result in over 3.3 million healthy life years lost globally each year [69]. In the area of LLM-aided mental health applications, the most frequently explored conditions include depression, suicide, and stress [2]. Eating disorders are understudied, though a recent study [70] that employed a closed-source LLM, ‘gpt-3.5-turbo-1106’, to analyze the psychosocial effects of ED on members of a Reddit forum. Yet, the study lacked evaluation by human or the use of multiple LLMs [48].
In this study, we evaluated the ensembling framework through two case studies of psychological textual data, focusing on eating disorders. We used human annotation to evaluate the labeling performances of the four open-source LLMs previously listed, as well as the ensembles of all possible combinations of them. For the public dataset of Reddit posts, we also used closed-source large-sized GPT-4o as a judge [48]. The Reddit posts, along with a human-annotated subset, are made public through the University of Virginia Dataverse [V3/NENELT_2025].
2.2.1 Case Study 1: Overlapping Topics Between ED and diet culture on Reddit
Social media has shown to promote diet culture, characterized by health myths about food and eating, as well as a moral hierarchy of bodies [71, 72]. Though diet culture lacks a uniform definition [71], it is defined in this study as a system of beliefs that equates thinness with health and oppresses people who do not align with standards of body shape, size, and weight [72]. Research shows that dieting and other weight-control methods significantly increase the risk of harmful eating practices associated with disordered eating or ED [73]. In this case study, we aimed to identify potentially overlapping topics between ED and general dieting online discussions to better understand the similarities between diet culture and ED. Extended from previous analyses of ED online discussions [74, 75], we identified 15 candidate topics that are likely prevalent in both dieting and ED communities, as listed in Table 2.2.1 (a).
{tableorg}
Descriptions of topic of interests in two empirical case studies 0pt Topic Short Name Description Binge eating binge Binge eating. Body dissatisfaction bodyhate Body dissatisfaction, feel bad about body image and appearance. Calorie count calorie Count calorie. Food cravings crave Craving for high calorie food or carbs. Depressed mood depressedmood Depressed mood, feeling depressed. ED recovery ed Eating disorders(ED) diagnosis or recovery, ED includes anorexia nervosa, anorexic, bulimia, bulimic, binge eating disorders, arfid, osfed, pica. Physical exercise exercise Physical exercise. Fear foods fearfood Fear certain foods. Fear of weight gain feargain Fear of body weight gain. Must related to body weight. Weight gain gain Body weight gain. Weight loss loss Body weight loss. High protein diet protein High protein diet, carbohydrate-reduced(low-carb) high-protein diet. Relationships relation Family and social relationships. Restriction restrict Restrict nutrition or calorie intake. Ideal body image idealbody Ideal body image including thinness, skinny body, low body fat and lean body mass. Case study 1: labeling overlapping topics between dieting and ED discussions among reddit posts 0pt Subcategories of Weight Stigma in Treatment Assumptions of health based on appearance or weight. Healthcare decisions based on weight. Treatment decision based on weight. Not sick enough, or not thin enought, or not low weight enough. Not taken seriously because of weight, Inadequate care based on weight. Diet promotion. Encouragement of weight loss. Reassurance of thinness. Reassurance that will not get fat in treatment. Negative attitudes, discrimination, or prejudice based on body weight or size. Weight blamed for health issues or concerns. Weight tied to personality characteristics. Case study 2: labeling weight stigma experienced by ed patients during treatment
We collected up to 1000 top-rated posts and up to 1000 newest posts per subreddit forum across three groups: dietary, fitness, and eating disorders, using Python package PRAW (7.7.1). Posts in dietary and fitness forums contain content of general dieting online discussions. In total, we collected 77,175 posts from 53,784 authors across 115 subreddit forums. Data descriptions and the designated searching keywords are shown in Table 1. For ED forums, we included 12 subreddits: ‘EatingDisorders’, ‘bulimia’, ‘intuitiveeating’, ‘AnorexiaNervosa’, ‘EDanonymemes’, ‘EDAnonymous’, ‘BingeEatingDisorder’, ‘fuckeatingdisorders’, ‘ARFID’, ‘eating_disorders’, ‘save_food’, ‘edsupport’. The data acquisition took place in February 2024.
As a public dataset, we conducted the evaluation in this case study using both human annotators and the large-size, closed-source GPT-4o. Firstly, for practicality, we randomly sampled 1,080 posts—10 posts per forum from 108 forums each containing at least 10 posts. One author labeled all 1,080 posts, and another author independently labeled a subset of 200 posts. The Gwet’s AC1 and Fleiss’ Kappa scores on the 200 posts labeled by the two coders were 0.972 (95% CI: 0.966 - 0.979) and 0.851 (95% CI: 0.817 - 0.886), respectively, indicating excellent agreement and reliability of the human annotation. Secondly, we used GPT-4o to label and rate all 77,175 posts. The Gwet’s AC1 and Fleiss’ Kappa with human annotation were 0.976 (95% CI: 0.974 - 0.979) and 0.885 (95% CI: 0.873 - 0.897), respectively, suggesting that large-sized LLMs can achieve close-to-human performance in these tasks, as previously recognized [76, 77].
Table 1: Case study 1 – Data Description
Forums Posts (Authors) Search by
ED 12 14853 (8878) eat disorder, anorexia nervosa, bulimia nervosa, binge eating disorder, ARFID, avoidant/restrictive food intake disorder, pica, other specified feeding or eating disorders, OSFED, unspecified feeding or eating disorders, orthorexia
Dietary 56 34367 (25459) diet, food, nutrition, meal, eat, fat, calorie, ingredient, menu
Fitness 47 27955 (20200) body shape, body image, body weight, thinness, exercise, fit, obesity
2.2.2 Case Study 2: ED Patients’ Experiences with Weight Stigma during Treatment
In the second case study, we used a protected clinical dataset of ED patients narratives about their experience with weight stigma during treatment. A total of 1,368 texts from patients treated for eating disorders were collected within a clinical setting. Two domain experts independently coded the dataset, determining the presence of weight stigma in patient narratives by identifying any of the subcategories listed in Table 2.2.1 (b). The Gwet’s AC1 and Fleiss’ Kappa between two coders’ annotation were 0.625 (95% CI: 0.583 - 0.667) and 0.541 (95% CI: 0.494 - 0.588), respectively. Due to the relatively low human inter-rater reliability, we used the subset of 1,080 texts where both coders agreed as the final human annotation to evaluate the labeling performance of LLMs and their ensembles.
In this case study, the LLMs’ labeling and scoring process resembled that of human labeling and was based on a codebook of weight stigma [78]. Each LLM was prompted to determine the presence of each subcategory, using the same detailed descriptions provided to human coders in the codebook. Subsequently, the label of weight stigma was calculated based on the presence of any subcategory, while the relevancy score was determined by the average relevancy scores of the present subcategories.
3 Results
3.1 Mitigating Heterogeneity through Relevancy Scoring
As shown in Table 3.2, in both case studies, the relevancy scores from LLMs indicated higher agreement, as measured by Gwet’s AC1 and Fleiss’ kappa, than the dichotomous labels returned by the LLMs. This result suggests that (1) the scoring method can effectively mitigate the heterogeneity in the labeling decisions made by multiple LLMs, and (2) the reasoning of LLMs exhibited higher similarity than their binary classifications.
3.2 Enhancing Accuracy through Ensembling
In case study 1, as shown in Figure 2, when using same-sized open-source LLMs to label overlapping topics between ED and dieting online discussions, there was large heterogeneity in their performances. Specifically, Qwen2.5-7B-Instruct exhibited low sensitivity but high precision, indicating it tended to be too conservative in predicting positive cases, yielding more false negatives. Conversely, Llama-3.1-8B-Instruct displayed high sensitivity but low precision, suggesting it was too generous in predicting positive cases, resulting in more false positives. Mistral-7B-Instruct-v0.3 had varying performances across different topics, while Vicuna-7b-v1.5 showed the lowest performance across all metrics.
However, the ensembles of LLMs overcame the shortcomings of individual LLMs and exhibited the highest performance in predicting human annotations across almost all topics. Firstly, the ensemble of four LLMs’ labels exhibited higher F1-score, sensitivity and precision than individual LLMs. Similarly, the PCA ensembles of scores showed the highest Area Under the Precision-Recall Curve (AUPRC) compared to individual LLMs. Specifically, the ensemble of four LLMs and the ensemble with outliers removed (final ensemble) showed significantly higher precision than all LLMs except for Qwen2.5-7B-Instruct (who had the lowest sensitivity), suggesting that ensembling can effectively identify and remove false positives. On the other hand, the 4-LLMs ensemble and the final ensemble exhibited higher sensitivity than all LLMs except for Llama-3.1-8B-Instruct (who had the lowest precision), suggesting that ensembling can effectively incorporate more true positives according to the agreement by multiple LLMs. Secondly, as shown in panel B, ensembling a larger number of LLMs resulted in higher performance gains, while using only two LLMs in the ensemble tended to be vulnerable to the heterogeneous performances of individual LLMs. Lastly, the accuracy of GPT-4o was generally higher than that of the selected open-source LLMs, as indicated by the dashed horizontal lines in Panel A. Using all 77,175 posts, the ensemble of LLMs demonstrated performance closer to GPT-4o compared to individual LLMs, as depicted in Figure 5. This result also suggests that the parameter size of an LLM should be considered an important factor in evaluating LLM-assisted tasks.
Figure 3 illustrates another comparison between the ensemble and individual LLMs, based on the distribution of their labels and scores across 77,175 posts. These are aggregated by the topical group of forums, with ’ED’ indicating eating disorder forums and ’dieting’ referring to dietary and fitness forums. Notably, the distribution of the average occurrence rate and relevancy score by the ensemble of LLMs most closely resembled those by GPT-4o and humans, compared to labels and scores by individual LLMs. There were large discrepancies in the distribution of labels and scores returned by individual LLMs.
In case study 2, as shown in Figure 4, LLMs exhibited varying capabilities in determining the presence of weight stigma in ED patients’ narratives. Qwen2.5-7B-Instruct was too conservative in labeling positive cases, resulting in low sensitivity and high precision or specificity, whereas Vicuna-7b-v1.5 was too generous, resulting in high sensitivity but low precision or specificity. Nevertheless, the ensemble of three or more LLMs achieved higher F1-score, precision, specificity and sensitivity than individual LLMs when classifying human annotations. Similarly, the ensembles of scores showed the highest AUPRC compared to individual LLMs. The removal of the detected outlier LLM, Qwen2.5-7B-Instruct, led to increased sensitivity but decreased precision and specificity. The overall F1-score remained the same as that of the ensemble of four LLMs. Note that specificity was included in the evaluation for this case study, due to the relatively balanced prevalence of human-annotated weight stigma–70% among 1080 posts.
{tableorg}
Evaluation of agreement on labeling and scoring1 conducted with four open-source, small-sized language models2. 0pt Gwet’s AC1 [95% CI] Fleiss’ kappa [95% CI] Topic Labels Scores Labels Scores binge 0.939 [0.937, 0.940] 0.924 [0.922, 0.925] 0.651 [0.644, 0.659] 0.702 [0.695, 0.709] bodyhate 0.827 [0.825, 0.830] 0.852 [0.850, 0.854] 0.451 [0.445, 0.457] 0.484 [0.477, 0.492] calorie 0.900 [0.898, 0.902] 0.881 [0.879, 0.883] 0.671 [0.665, 0.676] 0.722 [0.718, 0.727] crave 0.858 [0.856, 0.860] 0.884 [0.882, 0.886] 0.354 [0.347, 0.361] 0.408 [0.398, 0.418] depressedmood 0.774 [0.771, 0.777] 0.854 [0.852, 0.856] 0.321 [0.316, 0.325] 0.442 [0.435, 0.450] ed 0.902 [0.900, 0.904] 0.879 [0.877, 0.881] 0.689 [0.683, 0.694] 0.758 [0.753, 0.763] exercise 0.832 [0.829, 0.835] 0.781 [0.779, 0.784] 0.730 [0.726, 0.734] 0.647 [0.643, 0.651] fearfood 0.946 [0.944, 0.947] 0.946 [0.944, 0.947] 0.300 [0.291, 0.309] 0.354 [0.340, 0.369] feargain 0.904 [0.902, 0.905] 0.903 [0.901, 0.905] 0.377 [0.369, 0.385] 0.360 [0.351, 0.369] gain 0.880 [0.878, 0.882] 0.864 [0.862, 0.866] 0.563 [0.557, 0.568] 0.573 [0.567, 0.579] idealbody 0.790 [0.787, 0.792] 0.828 [0.825, 0.830] 0.246 [0.240, 0.251] 0.232 [0.225, 0.238] loss 0.751 [0.748, 0.754] 0.772 [0.769, 0.774] 0.546 [0.542, 0.551] 0.596 [0.591, 0.601] protein 0.796 [0.793, 0.799] 0.820 [0.818, 0.823] 0.452 [0.447, 0.458] 0.596 [0.589, 0.603] relation 0.338 [0.333, 0.342] 0.825 [0.823, 0.827] 0.057 [0.053, 0.061] 0.350 [0.343, 0.357] restrict 0.761 [0.757, 0.764] 0.793 [0.791, 0.796] 0.495 [0.490, 0.500] 0.526 [0.521, 0.532] Data 1: ED and Dieting Reddit Posts 0pt Gwet’s AC1 [95% CI] Fleiss’ kappa [95% CI] Topic Labels Scores Labels Scores weightstigma 0.601 [0.572, 0.629] 0.597 [0.575, 0.618] 0.187 [0.152, 0.222] 0.300 [0.269, 0.330] Data 2 (Protected): ED Patients Experiences
1 Embedding scores were calculated using all-mpnet-base-v2 and binned into 10 ordinal levels for agreement evaluation. 2 The four open- source, small-sized language models are Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct, Mistral-7B-Instruct-v0.3, and Vicuna-7b-v1.5.
Figure 2: Performance in predicting human labels of 15 topics related to eating disorders and dieting in Reddit posts, using four open-source, locally-deployable LLMs and all combinations of their ensembles. Panel A illustrates the performance of using relevancy scores or labels returned by LLMs and their ensembles to classify human annotations. The AUPRC metric evaluates relevancy scores, while F1-score, precision, and sensitivity evaluate performance based on labels. Different colors represent various LLMs and their ensembles. Performance metrics for GPT-4o are indicated by horizontal dashed lines. For ensembles of two and three LLMs, error bars indicate the minimum, maximum, and the 25th and 75th percentiles across all combinations of the respective ensemble size. Panel B illustrates the performance increases achieved by using ensembles of LLMs to classify human annotations, compared to the median performance of the individual LLMs included in each ensemble.
Figure 3: The average occurrence rates and average relevancy scores of 15 ED and dieting topics, as returned by LLMs and the ensemble, stratified by two forum groups: ED forums (blue) and dieting (dietary or fitness) forums (red). For comparison, the average occurrence rates labeled by GPT-4o and human annotators are highlighted with a gray background.
Figure 4: Performance in predicting human annotations of weight stigma experienced by ED patients during treatment, using four open-source, locally-deployable LLMs and all combinations of their ensembles. Panel A illustrates the performance of using relevancy scores or labels returned by LLMs and their ensembles to classify human annotations. The AUPRC metric evaluates relevancy scores, while F1-score, precision, and sensitivity evaluate performance based on labels. Different colors represent various LLMs and their ensembles. For ensembles of two and three LLMs, error bars indicate the minimum, maximum, and the 25th and 75th percentiles across all combinations of the respective ensemble size. Panel B illustrates the performance increases achieved by using ensembles of LLMs to classify human annotations, compared to the median performance of the individual LLMs included in each ensemble.
Figure 5: Performance in predicting GPT-4o labels of 15 topics related to eating disorders and dieting in Reddit posts, using four open-source, locally-deployable LLMs and all combinations of their ensembles. Panel A illustrates the performance of using relevancy scores or labels returned by LLMs and their ensembles to classify GPT-4o labels. The AUPRC metric evaluates relevancy scores, while F1-score, precision, and sensitivity evaluate performance based on labels. Different colors represent various LLMs and their ensembles. For ensembles of two and three LLMs, error bars indicate the minimum, maximum, and the 25th and 75th percentiles across all combinations of the respective ensemble size. Panel B illustrates the performance increases achieved by using ensembles of LLMs to classify GPT-4o annotations, compared to the median performance of the individual LLMs included in each ensemble.
4 Discussion
In psychology research, annotating granular topics, such as pathological behaviors and mentalities in free-form texts like patient narratives, is typically labor-intensive and time-consuming. Without user consent for external use, LLM-assisted labeling techniques that rely on closed-source models through APIs can hardly be applied to such free-text data. These constraints also limit the use of large-sized LLMs ($>$100 billion parameters) because they cannot be deployed locally in most research settings.
In this study, we proposed an ensembling framework that leverages the heterogeneity of diverse, open-source, locally-deployable LLMs for topic labeling and relevancy scoring on free-text data. This ensembling is guided by a novel relevancy scoring methodology that utilizes embedding distances between the topic description and LLMs’ reasoning. The topic labeling and scoring by multiple LLMs mirror the human labeling process performed by several human coders. The main idea of ensembling labels from multiple LLMs is to eliminate potential false positives from the union of positive labels returned by multiple LLMs, which inherits all false positives from each LLM. Conversely, the intersection of positive labels, based on over-half agreement among LLMs, is likely to exclude more true positives. Thus, label ensembling seeks a balance between the union and the intersection, guided by the ensemble of relevancy scores.
The framework was not designed to replace human evaluation, but rather to alleviate the burden of the human labeling process by providing a preliminary label and a relevancy score for each text. The final labeling decision, typically made by domain experts, can reference these preliminary labels, focusing particularly on positive labels with low relevancy scores (potential false positives) and negative labels with positive scores (potential false negatives).
In the experimentation, we deployed four small-sized (7-8 billion parameters) LLMs locally: Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct, Mistral-7B-Instruct-v0.3, Vicuna-7b-v1.5, and evaluated the framework through two case studies on human-annotated psychological textual data, with a focus on eating disorders (ED). The first study analyzed overlapping topics between ED and dieting across 77,175 public Reddit posts, while the second documented ED patients’ experiences with weight stigma during treatment in a free-text dataset of 1,080 patient narratives. We found that:
  1. 1.
    Evaluated by human annotations, there is heterogeneity among the selected same-sized LLMs when conducting the same topic labeling tasks, with some exhibiting low sensitivity but high precision, and others high sensitivity but low precision. Such heterogeneity reflects the varied data, model architectures, and preprocessing steps each LLM was exposed to during training, as well as constraints imposed by the size of their parameters.
  2. 2.
    The ensemble of LLMs effectively leveraged the heterogeneity of the LLMs, achieving the highest accuracy in predicting human annotations. In the first case study, the ensemble of LLMs also demonstrated the highest accuracy in predicting closed-source large-sized GPT-4o’s labels. The improvement in accuracy was evidenced by superior F1-score, superior precision, and sensitivity compared to individual LLMs.
  3. 3.
    The ensemble of a larger number of LLMs demonstrated more reliable performance improvements and a more optimal precision-recall trade-off. In contrast, ensembles of only two LLMs appeared more vulnerable to the variability in individual LLM performance.
  4. 4.
    The relevancy scores across LLMs showed greater agreement than dichotomous labels, as measured by Gwet’s AC1 and Fleiss’ kappa. This suggests the proposed relevancy scoring method can effectively mitigate the variability in LLMs’ labeling.
Nevertheless, given the proposed ensembling framework, potential extensions at each stage warrant further investigation. First, while this study focused on ensembling open-source, deployable LLMs for free-text data, the framework could be expanded to include any textual data and larger-sized LLMs. For example, we controlled for the parameter size of LLMs to be between 7-8 billion parameters in our experiments. Larger-sized open-source LLMs, such as Llama-3.1-70B-Instruct and Llama-3.1-405B [34], may warrant examination in future studies that have access to more computational resources. Second, techniques such as few-shot prompting [54], chain-of-thought [55], and other prompt engineering techniques [56, 57] can be employed as alternatives for the simple prompting in this experimentation. The potential of performance improvement by advanced prompting may result in improved accuracy for both individual LLMs and the ensemble of them. Third, the relevancy scoring utilizes the embedding distance between topic description and LLM-returned related phrases. Besides the pre-trained Sentence-BERT model ‘all-mpnet-base-v2,’ other pre-trained embedding models, whether encoder-decoder or decoder-only, may serve as potential alternatives. Lastly, an intriguing direction for future study is to fine-tune individual LLMs based on the ensemble of labels and scores for a given topic. This approach may provide individual models with unseen information and bridge their capacity to perform inference tasks in specific psychological contexts.